### v1

# Assessing the Carbon Footprint of Virtual Meetings: A Quantitative Analysis of Camera Usage

Abstract
This paper analyzes the carbon emissions related to data consumption during video calls, focusing on the impact of having the camera on versus off. Addresses the energy efficiency and carbon footprint of digital communication tools. The study is used to quantify the real reduction in environmental impact claimed in several articles when people choose to turn off their camera during meetings. The experiment was carried out using a 4G connection via a cell phone to understand the varying data transfer associated with videos. The findings indicate that turning the camera off can halve data consumption therefore carbon emissions, particularly on mobile networks, and conclude with recommendations to optimize data usage and reduce environmental impact during calls.
I Introduction
The Information and Communication Technology (ICT) sector is responsible for approximately 2 to 4% of global greenhouse gas (GHG) emissions [Freitag2021]. With the advent of generative AI, reputable institutions project a significant increase in these emissions in the near future [Iea2025]. To mitigate the environmental impact of this growth, both IT developers and end-users are increasingly adopting responsible practices.
A prevalent assertion suggests that deactivating cameras during virtual conferences may mitigate GHG emissions. Few serious studies have been conducted to determine to what extent the environmental impact is reduced when this best practice is used. The most popular and serious study on this subject is highlighting a reduction in 96% environmental impact by turning off cameras during virtual meetings [Obringer2021]. However, while the findings are promising, they are largely based on estimates and modeling rather than direct experimental data. To obtain more accurate and actionable insights, further studies that include real-world experiments are necessary.
This report quantifies and analyzes the data consumption during video calls, focusing on the impact of having the camera on versus off. The analysis includes a preliminary statistical test and recommendations based on the findings.
The consumption of 4G data is closely linked to carbon dioxide equivalent (CO2e) emissions. The energy consumption of mobile networks such as 4G varies significantly according to the volume of data transferred. According to some estimates, transferring one gigabyte (GB) of data on 4G corresponds to a consumption of electricity of 0.6 kWh [Arcep2019] which contributes directly to CO2 emissions associated with the use of these technologies.
In contrast, fixed networks such as fiber optics and ADSL have a different energy profile. According to a statement by the ARCEP French organization, the energy consumption of fixed networks depends relatively little on how they are used. For example, fiber optics consume an average of 0.5 W, with little variation, regardless of the amount of data transmitted [Arcep2019]. The energy efficiency of fiber networks is due to the physical advantage of light that does not require active network elements, as well as the high data rates allowed by the technology. The largest consumer of electricity in fixed networks is the customer premises equipment (CPE), including a modem router unit and a possible termination of the optical network [Prysmian2021]. This means that even with intensive use, the impact on electricity consumption remains marginal.
Although fairly stable, data on electricity consumption related to data transfer over fixed networks have been published in the literature.
TABLE I: Energy consumption of different networks.
Network kWh/GB Country Year
Fix (ADSL) 0.0002 Germany -
Fix (Fiber) 0.00002 Germany -
Fix 0.1 Finland 2020
Fix 0.03 Europe 2020
Fix 0.05 Finland 2022
Fix (Fiber) 0.04 USA -
I presents data on energy consumption for fixed (wired) network communications, measured in kilowatt-hours per gigabyte (kWh/GB).
For fixed ADSL networks, [Prysmian2021] report an energy consumption of approximately 0.0002 kWh/GB in Germany. Similarly, fixed fiber networks in Germany also show an energy consumption of 0.00002 kWh/GB as reported by the same source [Prysmian2021]. Data for the USA from [Thundersaid2024] shows that fixed fiber networks consume 0.04 kWh/GB. In Finland, [Pihkola2018] indicate a higher consumption of 0.1 kWh/GB for fixed networks. Data for Europe, provided by [EuropeanCommission2020], is showing that fixed networks consume 0.03 kWh/GB. Additionally, [Traficom2022] reports a consumption of 0.05 kWh/GB for fixed networks in Finland.
Although data has been found for electricity consumption per amount of data transferred on fixed networks ranging from 0.00002 to 0.1 kWh/GB, we’re going to retain ARCEP’s assumption that the amount of electricity consumed varies little according to the amount of data transferred for the continuation of the study, while bearing in mind that it still has a small impact.
This paper is structured as follows: Section II describes the methodology used for the study. Section III presents the data overview and preliminary analysis. Section IV discusses the interpretation of the results. Section V provides recommendations based on the findings. Finally, Section VI concludes the paper and discusses future work.
II Methodology
The meetings were conducted using a 4G connection via a cell phone. At the start of each meeting, mobile data consumption for the month was recorded using the Red by SFR French operator application [Sfr]. The following data were then noted: the start time, the number of participants, the number of cameras switched on, as well as whether a participant was sharing his or her screen, and the content shared. Each time one of these variables changed during the meeting, the time of the change and the variable concerned were recorded.
At the end of each meeting, mobile data consumption was again recorded, enabling the difference from the initial measurement to be calculated. This gave the specific data consumption for each meeting. In addition, all the variables noted were summed together to analyze their relationship with data consumption.
A total of 9 meetings were held, during between 30 to 60 minutes. Meetings were classified as “camera on” when the cameras were on for almost the entire duration, and as “camera off” when the cameras were off for the majority of the meeting.
The hardware used in the experiment consisted of an HP Elitebook laptop with an Intel Core i7 processor [Hp] using Windows 10 operating system [Windows10]. For videoconferencing meetings, the software used was Microsoft Teams [Teams]. No other windows were open during the majority of the meetings, which means that data consumption is exclusively linked to the meeting itself. On the cell phone side, no applications were launched for the duration of the meetings. Thus, data consumption can also be attributed solely to the meeting.
III Data Overview and Preliminary Analysis
The dataset includes details on call participants, data consumption, call duration, and various screen sharing activities. The key metrics analyzed are:
  • Data consumption per minute with the camera on.
  • Data consumption per minute with the camera off.
The data electricity consumption factors come from different sources and can be used to define a range of kWh consumed:
TABLE II: Energy consumption of different networks.
Network kWh/GB Country Year
Mobile 0.14 Europe 2020
4G 0.17 Finland 2020
4G 0.26 Belgium 2020
Mobile (3G, 4G, 5G) 0.12 Finland 2022
II presents a comparative analysis of energy consumption in kilowatt-hours per gigabyte (kWh/GB) for video calls using various mobile networks across different countries and years.
[EuropeanCommission2020] report an energy consumption of 0.14 kWh/GB in Europe for the year 2020 for mobile network, not precising the technologie used. In Finland, [Nokia2025] indicates that the 4G network consumes 0.17 kWh/GB. [Golard2020] present data for Belgium, where the 4G network consumes 0.26 kWh/GB in 2020. Additionally, [Traficom2022] provides a comprehensive analysis for Finland, showing that mobile networks, including 3G, 4G, and 5G, consume 0.12 kWh/GB in 2022.
Even though it has been stated that the study is based on the ARCEP [Arcep2019] assumption that the amount of data transferred does not vary the amount of electricity consumed on the fixed network, it can still be noted that electricity consumption in kWh per GB is 2.6 to 6 000 times higher on the mobile network than the data found for the fixed network.
Using a French ADEME Base Empreinte 2013 carbon emission factor of 79.1 g of CO2e/kWh [Baseempreinte], electricity consumption contributes to 9.492 to 20.566 kg of CO2e emissions associated with the use of these mobile networks.
Assuming 1 GB = 1024 MB, this leads to emission factors values from 9.270 to 20.084 mg of CO2e/MB.
We did not conduct a t-test to compare the average data consumption per minute with the camera on versus off, as the descriptive statistic is sufficient:
  • Average data consumption per minute with camera on: 18.07 MB
  • Average data consumption per minute with camera off: 8.76 MB
There is approximately a 2x increase in data consumption when the camera is on compared to when it is off during meetings.
IV Interpretation
The previous analysis indicates a statistically significant difference in data consumption between having the camera on and off during video calls. However, this analysis is preliminary and does not account for screen sharing activities or instances where the camera was switched on or off during the call.
Converting these quantities of mobile data into CO2e using the French emission factor provided by ADEME, the additional carbon emissions of a one-hour video call with a camera to one without can be compared with several use cases.
TABLE III: CO2 emissions of different activities.
Activity 1 minute 1 hour 1 unit
in g CO2e
Call camera-on 0.17 – 0.36 10.2 – 21.6
Call camera-off 0.08 - 0.18 4.8 – 10.8
Watching Netflix 0.6 36
GPT-3.5 request 0.6
Hot shower 18 1 080 162
III provides a detailed comparison of CO2e emissions associated with various activities, measured in grams. For video calls, having the camera on results in emissions ranging from 0.17 to 0.36 grams of CO2e per minute, which translates to 10.2 to 21.6 grams over an hour. Turning the camera off reduces these emissions to 0.08 to 0.18 grams per minute, or 4.8 to 10.8 grams per hour. Watching Netflix is shown to emit 0.6 grams of CO2e per minute, accumulating to 36 grams over an hour. A single request using GPT-3.5 generates 0.6 grams of CO2e. Additionally, taking a hot shower emits 18 grams of CO2e per minute [Carbonfootprintmodelingtool], 1,080 grams per hour, and 162 grams per unit.
Based on III, a one-hour video call with a camera on is equivalent to about a 1-minute hot shower or a 30-minutes netflix session.
V Recommendations
It is advisable to systematically use fixed networks such as Ethernet or Wi-Fi whenever possible. The quantity of data transferred does not significantly impact electricity consumption on these networks. For calls with a small number of participants, turning the camera on can foster social connection and happiness [Sherman2013], which in turn can enhance motivation and productivity. However, during large meetings or webinars where active participation is not required, it is recommended to turn the camera off. Even though the impact is not significant, the quantity of data transferred has a small effect on electricity consumption on these networks.
When using mobile networks such as 3G, 4G, or 5G, it is best to turn the camera off unless it is essential to see each other. Even in environments where visual interaction is important, it is crucial to consider the impact on data consumption and electricity usage.
VI Conclusion and Future Work
The preliminary analysis indicates that turning the camera off during video calls can halve data consumption and significantly reduce carbon footprint, particularly on mobile networks. While further analysis is needed to account for screen sharing and dynamic camera usage, the current recommendations emphasize the use of fixed networks and judicious camera usage to optimize data consumption and electricity usage, therefore carbon footprint.
To provide a more comprehensive analysis, it is recommended to incorporate screen sharing data by analyzing the impact of different types of screen sharing, such as PowerPoint, Excel, Internet, and Code, on data consumption. This involves determining if screen sharing activities significantly affect data consumption when the camera is on versus off. Additionally, accounting for instances where the camera is switched on or off during the call will allow for a more detailed analysis to understand the dynamic impact of camera usage on data consumption. Furthermore, expanding the dataset by collecting additional data will strengthen the statistical significance of the results.
VII Acknowledgments
We would like to thank Boris RUF for his careful proof-reading and invaluable advice.